ANN'SAGE IS 7/10TH OF BILL'S AGE. IN 4 YEARS ANN'S AGE WILL BE 8/11THS OF BILL'S AGE. HOW OLD IS BILL?
Found 2 solutions by alicealc, lazi:
Answer by alicealc(293)

Ann's age = A
Bill's age = B

A = 7/10 * B
In 4 years:
A + 4 = 8/11 * (B + 4)
7/10 * B + 4 = 8/11 * B + 32/11
7/10 * B - 8/11 * B = 32/11 - 4
(77/110 - 80/110) * B = 32/11 - 44/11
(-3/110) * B = -12/11
B = -12/11 * -110/3
B = 40
A = 7/10 * 40 = 28

so,Bill is 40 years old now



Answer by lazi(15)

a=Ann's age
b= Bill's age
a= (7/10)b = (7b)/10
in four years means
a + 4 = (8/11)(b+4)
now we know what a is in terms of b, we just plug it into the formula above
(7b)/10 + 4 = (8b)/11 + 32/11
solve for b
(7b)/10 + 4 = (8b)/11 + 32/11
-(7b)/10 on both side
-32/11 on both side
_____________________________
12/11 = (3/110)b
times 110/3 on both side
_________________
40 = b
Bill's age is 40.